OpenStudy (anonymous):

2x^2− 4 = 8 find the smaller root

OpenStudy (zepp):

You can solve it by using the quadratic formula, set one side to 0

OpenStudy (zepp):

Or by doing 2x^2 = 8 + 4 2x^2 = 12 x^2 = 12/2 x^2 = 6 \[x=\pm \sqrt{6}\] Smallest root will be \[x=-\sqrt{6}\]